# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ Appendix B GPARV for Laplace Mechanism
$\omega(y)=\inf_{x\in\{0,1\}}\frac{\mathcal{R}(x)(y)}{\gamma}=\begin{cases}\frac{\varepsilon_{0}}{2\gamma}\exp\left(\varepsilon_{0}(y-1)\right),&y\leq 0.5,\\ \frac{\varepsilon_{0}}{2\gamma}\exp\left(-\varepsilon_{0}y\right),&y>0.5,\end{cases}$
@@ Appendix C Proof of Theorem 12
The proof of this theorem follows similar lines as the proof of Lemma 12 in [BBGN19].
@@
[$\mathbb{P}[\{Y_{1}^{b}\}\cup\mathbf{W}_{n-1}=Y]=\frac{1}{n!}\sum_{\sigma}\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right],$
@@
10-RR, low $\varepsilon_{0}$
10-RR, large $\varepsilon_{0}$
@@
Figure 2: Experimental Results: RR, BLH and RAPPOR
@@
$\frac{1}{2}$+$\frac{1}{2}$, low $\varepsilon_{0}$
$\frac{1}{2}$+$\frac{1}{2}$, large $\varepsilon_{0}$
@@
Figure 6: 10-RR under Poisson Subsampling (with sampling rate $r$) and Shuffling

### v6

@@ Appendix B GPARV for Laplace Mechanism
$\omega(y)=\inf_{x\in[0,1]}\frac{\mathcal{R}(x)(y)}{\gamma}=\begin{cases}\frac{\varepsilon_{0}}{2\gamma}\exp\left(\varepsilon_{0}(y-1)\right),&y\leq 0.5,\\ \frac{\varepsilon_{0}}{2\gamma}\exp\left(-\varepsilon_{0}y\right),&y>0.5,\end{cases}$
@@ Appendix C Proof of Theorem 12
The proof of this theorem follows similar lines as the proof of Lemma 12 in [Balle2019].
@@
$\mathbb{P}[\{Y_{1}^{b}\}\cup\mathbf{W}_{n-1}=Y]=\frac{1}{n!}\sum_{\sigma}\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right],$
@@
10-$R$, low $\varepsilon_{0}$
10-$R$, large $\varepsilon_{0}$
@@
Figure 2: Experimental Results: $R$, BLH and RAPPOR
@@
[…]
@@
Table 4: Frequency estimation under 10-$R$ ($n=1000,\delta=10^{-6}$): each entry reports $(\varepsilon_{0},\ \ell_{2}\text{ error})$ required to achieve target privacy level $\varepsilon$.
$\varepsilon$ 0.01 0.05 0.10 0.20 0.50 1.00
Blanket (Bennett) (0.170, 1.598)(0.170,\ 1.598) (0.620, 0.366)(0.620,\ 0.366) (1.000, 0.199)(1.000,\ 0.199) (1.510, 0.108)(1.510,\ 0.108) (2.370, 0.051)(2.370,\ 0.051) (3.060, 0.033)(3.060,\ 0.033)
Our FFT (0.210, 1.285)(0.210,\ 1.285) (0.730, 0.295)(0.730,\ 0.295) (1.150, 0.163)(1.150,\ 0.163) (1.700, 0.089)(1.700,\ 0.089) (2.650, 0.042)(2.650,\ 0.042) (3.510, 0.025)(3.510,\ 0.025)
@@
Figure 6: Upper Bounds for 10-$R$ under Poisson Subsampling and Shuffling